# Neutrinos from explosive transients at the dawn of multi-messenger astronomy

### v1

@@ 3.1 Non-thermal neutrino emission from transient astrophysical sources
The IceCube Neutrino Observatory has detected neutrinos of astrophysical origin with $\rm{TeV}$$\rm{PeV}$ energy since 2013. A handful of statistically significant neutrino-electromagnetic associations has been found, e.g. in coincidence with the blazar event TXS 0506$+$056, the active galaxy NGC 1068, and our own Galaxy 142; 3. Additionally the number of possible multi-messenger associations steadily increases 143. A very intriguing opportunity would stem from the detection of high-energy neutrinos from transients originating from collapsing massive stars and neutron-star mergers. In fact, such neutrinos can shed light on the mechanism powering the observed electromagnetic emission as well as the properties of the source and its surrounding medium 144; 145.
@@ 3.1.1 Collapsing massive stars and emerging transient sources
In addition to thermal neutrinos with energy of $\mathcal{O}(10)$ MeV, the explosion resulting from the collapse of a massive star can also lead to the production of non-thermal neutrinos with higher energy (in the TeV–PeV range). Such neutrinos are produced because of the interactions of accelerated protons with other protons or photons (hadronic or photo-hadronic interactions) 146; 147. A summary of the particle acceleration sites, leading to non-thermal neutrino production for high-energy transients stemming from collapsing massive stars, is provided in the left panel of Fig. 3.
@@
The supernova ejecta expand at 333310×$10\times 10^{3}$ km/s, and diffusive shock acceleration can take place in the proximity of the shock. In shock-powered supernovae, high-energy neutrinos could be produced because of the interaction of shock-accelerated protons with protons from the dense circumstellar medium (gamma-ray emission should be expected as well), or with the photon spectrum at the external shock (albeit the related neutrino contribution is subleading for non-relativistic or mildly-relativistic shocks)–see the left panel of Fig. 3. Such non-thermal particle emission is efficient for supernovae of Type IIn, IIp, or superluminous supernovae (cf. Table 1 and Fig. 4), whose circumstellar medium is expected to be especially dense because of the collapsing star ejecting its outermost layers 37; 44. High-energy neutrinos from these sources have not been detected yet, with the IceCube Neutrino Observatory providing upper limits on the diffuse emission from (shock-powered) supernovae 148. However, the detection of these neutrinos would be important to explore the structure
@@
and properties of the circumstellar medium as well as the ejecta. Interacting supernovae can be among the main contributors to the diffuse neutrino flux between $10$ and $10^{3}$ TeV, without overshooting the isotropic gamma-ray background observed by Fermi-LAT 37; 152.
If the central compact object is a magnetar, we expect neutrino emission linked to its spin down 153. In this case, the protons accelerated in the magnetar wind can undergo hadronic or photo-hadronic interactions with the thermal optical photons and the non-thermal X-rays produced in the ejecta forming a wind nebula, as shown in Fig. 3.
A jet can be harbored in stripped supernovae (of Type Ib/c), leading to gamma-ray bursts (cf. Table 1 and Fig. 3). The jet may be successful or not, according to its energy and the density of the stellar envelope. Below the photosphere, the mixing between the jet and the cocoon loads the jet with baryons, inhibiting subphotospheric collisionless shocks 154. However, if the jet is even mildly magnetized, magnetic reconnection and subshocks may favor particle acceleration and therefore neutrino production in the GeV–TeV range 155; 156. In the optically thin region of the relativistic jet, observed as a gamma-ray burst, the accelerated protons can interact with non-thermal photons to produce high-energy neutrinos 149; 53, as displayed in Fig. 4. Notably the efficiency of particle acceleration depends on whether heavier nuclei can be entrained in the jet, with neutrino emission becoming less efficient 157; 158; 159. To date, neutrinos emitted from gamma-ray bursts have not yet been detected, these findings are however compatible with theoretical estimations 160; 149.
The prospects of detecting neutrinos from newly discovered source classes potentially linked to collapsing massive stars are also very exciting. An example is provided by luminous fast blue optical transients (LFBOTs, see Table 1), such as AT2018cow 161. The origin of these transients is not yet known, but they are expected to be powered by a compact object, launching an asymmetric and fast outflow 162. It has been proposed that luminous fast blue optical transients may stem from collapsing massive stars stripped of their hydrogen envelope and harboring a choked jet surrounded by an inflated cocoon 163; the related neutrino fluence is plotted in Fig. 4 for reference. Alternatively, the luminous fast blue optical transient electromagnetic emission could come from the delayed merger of a black hole with a Wolf-Rayet star 164. Neutrinos could be produced because of the interaction between the outflow and the circumstellar medium 151 or in the magnetar wind 153. The detection of neutrinos would be crucial to breaking the degeneracies plaguing the electromagnetic data (which currently do not allow for a robust assessment of the source properties), allowing to pinpoint the mechanism powering the observed electromagnetic emission 151.
@@ 3.1.2 Neutron-star mergers
Following up the merger of two neutron stars, if the remnant compact object is a magnetar, its wind can interact with the ejecta, with protons being accelerated at the magnetosphere or the termination shock of the wind, see the right panel of Fig. 3. While particle acceleration is not efficient at the termination shock, neutrino emission from the magnetar could be detected in connection with gravitational waves 165; however, recent work suggests that the detection of such neutrinos may be more challenging than initially expected 166.
The non-detection of neutrinos from GRB 170817A 6, observed in connection with gravitational waves, is consistent with the fact that the relativistic jet was off-axis with respect to the observer 154; 57 (cf. Fig. 4 for an estimation of the muon neutrino fluence expected from the prompt emission). As the jet propagates within the ejecta, particle acceleration can take place at internal shocks inside the ejecta, leading to a neutrino signal within the detection reach of the IceCube Neutrino Telescope 47. Table 1 provides a summary of the energetics and the rate of the transients linked to the coalescence of neutron stars, with the related acceleration sites being summarized in the right panel of Fig. 3.
@@ 3.2 Best detection strategies for upcoming high-energy multi-messenger searches
The detection (or lack thereof) of high-energy neutrinos from transients stemming from collapsing massive stars or neutron-star mergers allows to gain crucial insight about the source properties, the behavior of plasma in highly magnetized media, as well as the efficiency of particle acceleration. Of course, the information provided by high-energy neutrinos should be complemented by the one carried by photons across wavebands and gravitational waves. Notably, detecting non-thermal neutrinos is especially crucial for electromagnetically dark sources, such as choked jets. In this case, we may expect a correlation of high-energy neutrinos with gravitational waves and $\mathcal{O}(10)$ MeV neutrinos (or interactions among such neutrinos 167) produced in the source core.
@@
Within the next decade, the number of electromagnetic detections of transients is expected to grow swiftly thanks to wide field, high-cadence surveys, such as the soon-to-be operative Vera C. Rubin Observatory 168 and ULTRASAT 169, the Nancy Grace Roman Space Telescope 170 as well as the Cherenkov Telescope Array (CTA) 171. These surveys, among others, will complement the pioneering findings of the Zwicky Transient Facility (ZTF) 172, the Panoramic Survey Telescope and Rapid Response System 1 (Pan-STARRS1) 173, the Swift-Ultra-Violet/Optical Telescope 174, and the James Webb Space Telescope 175. At the same time, the next generation gravitational wave detectors, e.g. the Cosmic Explorer 176 and Einstein Telescope 177, are expected to largely expand the observational horizon explored by LIGO, Virgo, and KAGRA 178. For what concerns the detection of high-energy neutrino counterparts, the current observational landscape is dominated by the IceCube Neutrino Telescope 3 and the under-construction Cubic Kilometre Neutrino Telescope (KM3NeT) 179. The planned IceCube-Gen2 180, the Giant Radio Array for Neutrino Detection (GRAND200k) 181, and the Pacific Ocean Neutrino Experiment (P-ONE) 182, among several other neutrino detectors, are expected to strengthen and expand our detection opportunities (cf. Table 1 of 142 for an overview of existing and planned high-energy neutrino telescopes) 3.
@@
To investigate the source properties and test our understanding of particle acceleration physics, it is however necessary to facilitate coordination as well as data exchange through multi-messenger programs. Several joint programs are already operative, for example ZTF, Pan-STARRS1, the Dark Energy Survey (DES), and the All-Sky Automated Survey for SuperNovae (ASAS-SN) carry out target-of-opportunity searches for optical counterparts of neutrino events and vice versa 183; 184; 185; 186; similar collaborations exist for neutrino and gamma-ray joint searches, involving Fermi-LAT and the Imaging Atmospheric Cherenkov Telescopes (IACTs) 187; 188. To take full advantage of upcoming joint high-energy neutrino and electromagnetic detections, it is crucial to strengthen correlations between neutrinos, the radio waveband (likely carrying imprints of interactions with the circumstellar medium), and the X-ray one (sensitive to the activity of the central engine) to break the degeneracies intrinsic to the optical-ultraviolet-infrared band 44; 144; 152. The time window adopted to look for neutrinos should not be chosen arbitrarily; in fact, we expect that the neutrino signal can only be tightly linked to the waveband whose photons are produced in the same source region.
@@ 4 Neutrinos and explosive transients: future challenges and opportunities
The growing number of likely multi-messenger detections of transient astrophysical sources, as well as the major developments taking place in the field of neutrino telescopes and gravitational wave detectors, jointly with the upcoming wide field and high-cadence surveys revolutionizing electromagnetic observations of transients, call for an improved understanding of the source physics and related particle production. Only by advancing the theoretical modeling, we would be able to take advantage of the upcoming plethora of data. On the one hand, it is essential to make progress in the modeling of neutrino flavor conversion in dense media. One challenge is that neutrino flavor conversion is not taken into account in (magneto)hydrodynamic simulations of the source. Hence, the impact of the feedback of flavor conversion physics on the source and the nucleosynthesis is yet to be understood. On the other hand, it is crucial to swiftly gain ground in our understanding of particle acceleration. Recent work highlights the importance of coupling particle transport consistently with general relativistic magnetohydrodynamics and particle-in-cell simulations to reliably assess the regions (and the efficiency) of particle acceleration and forecast the multi-messenger signals. For emerging source classes, neutrinos will be important to breaking the degeneracies affecting the electromagnetic signal, pinpointing the physics mechanisms powering such sources.
@@
The unprecedented number of real-time alerts expected to become available to telescopes involved in multi-messenger searches calls for a rapid development of the cyberinfrastructure needed to coordinate, cross-correlate, and analyze the large amount of data provided by such telescopes. In this respect, optimization of the communication infrastructure adopted to broadcast alerts is urgently needed. The Scalable Cyberinfrastructure to support Multi-Messenger Astrophysics (SCiMMA) 189 and the Time-domain Astronomy Coordination Hub (TACH) 190 are examples of projects in this direction.
Expanding the detection horizon beyond a single source, the diffuse emission of neutrinos from specific source classes can inform us on the source population properties. Besides complementing electromagnetic measurements, diffuse neutrino data could provide input on the sources that are dim or dark electromagnetically. However, to be able to extract details from the neutrino signal and predict the diffuse signal reliably, we should proceed on the modeling of stellar evolution, as well as better understand the role of a binary companion in the stellar evolution and the cosmic rate.

### v2

@@ 3.1 Non-thermal neutrino emission from transient astrophysical sources
The IceCube Neutrino Observatory has detected neutrinos of astrophysical origin with $\rm{TeV}$$\rm{PeV}$ energy since 2013. A handful of statistically significant neutrino-electromagnetic associations has been found, e.g. in coincidence with the blazar event TXS 0506$+$056, the active galaxy NGC 1068, and our own Galaxy 3; 4; 5. The potential for multi-messenger follow-up in real-time using neutrino alert is steadily improving 152; at the same time, the overall number of detected high-energy neutrino events increases 153, offering additional opportunities to search for multi-messenger associations. A very intriguing opportunity would stem from the detection of high-energy neutrinos from transients originating from collapsing massive stars and neutron-star mergers. In fact, such neutrinos can shed light on the mechanism powering the observed electromagnetic emission as well as the properties of the source and its surrounding medium 154; 155.
@@ 3.1.1 Collapsing massive stars and emerging transient sources
In addition to thermal neutrinos with energy of $\mathcal{O}(10)$ MeV, the explosion resulting from the collapse of a massive star can also lead to the production of non-thermal neutrinos with higher energy (in the TeV–PeV range). Such neutrinos are produced because of the interactions of accelerated protons with other protons or photons (hadronic or photo-hadronic interactions) 156; 157. A summary of the particle acceleration sites, leading to non-thermal neutrino production for high-energy transients stemming from collapsing massive stars, is provided in the left panel of Fig. 3.
@@
The supernova ejecta expand at 333310×$10\times 10^{3}$ km/s, and diffusive shock acceleration can take place in the proximity of the shock. In shock-powered supernovae, high-energy neutrinos could be produced because of the interaction of shock-accelerated protons with protons from the dense circumstellar medium (gamma-ray emission should be expected as well), or with the photon spectrum at the external shock (albeit the related neutrino contribution is subleading for non-relativistic or mildly-relativistic shocks)–see the left panel of Fig. 3. Such non-thermal particle emission is efficient for supernovae of Type IIn, IIp, or superluminous supernovae (cf. Table 1 and Fig. 4), whose circumstellar medium is expected to be especially dense because of mass-loss ejection from the progenitor star 158; 44; 51. High-energy neutrinos from these sources have not been detected yet, with the IceCube Neutrino Observatory providing upper limits on the diffuse emission from (shock-powered) supernovae 159. However, the detection of these neutrinos would be important to explore the structure
@@
and properties of the circumstellar medium as well as the ejecta. Interacting supernovae can be among the main contributors to the diffuse neutrino flux between $10$ and $10^{3}$ TeV, without overshooting the isotropic gamma-ray background observed by Fermi-LAT 44; 163.
If the central compact object is a magnetar, we expect neutrino emission linked to its spin down 164. In this case, the protons accelerated in the magnetar wind can undergo hadronic or photo-hadronic interactions with the thermal optical photons and the non-thermal X-rays produced in the ejecta forming a wind nebula, as shown in Fig. 3.
A jet can be harbored in stripped supernovae (of Type Ib/c), leading to gamma-ray bursts (cf. Table 1 and Fig. 3). The jet may be successful or not, according to its energy and the density of the stellar envelope. Below the photosphere, the mixing between the jet and the cocoon loads the jet with baryons, inhibiting subphotospheric collisionless shocks 165. However, if the jet is even mildly magnetized, magnetic reconnection and subshocks may favor particle acceleration and therefore neutrino production in the GeV–TeV range 166; 167. In the optically thin region of the relativistic jet, observed as a gamma-ray burst, the accelerated protons can interact with non-thermal photons to produce high-energy neutrinos 160; 60, as displayed in Fig. 4. Notably the efficiency of particle acceleration depends on whether heavier nuclei can be entrained in the jet, with neutrino emission becoming less efficient 168; 169; 170. To date, neutrinos emitted from gamma-ray bursts have not yet been detected, these findings are however compatible with theoretical estimations 171; 160.
The prospects of detecting neutrinos from newly discovered source classes potentially linked to collapsing massive stars are also very exciting. An example is provided by luminous fast blue optical transients (LFBOTs, see Table 1), such as AT2018cow 172. The origin of these transients is not yet known, but they are expected to be powered by a compact object, launching an asymmetric and fast outflow 173. It has been proposed that luminous fast blue optical transients may stem from collapsing massive stars stripped of their hydrogen envelope and harboring a choked jet surrounded by an inflated cocoon 174; the related neutrino fluence is plotted in Fig. 4 for reference. Alternatively, the luminous fast blue optical transient electromagnetic emission could come from the delayed merger of a black hole with a Wolf-Rayet star 25. Neutrinos could be produced because of the interaction between the outflow and the circumstellar medium 162 or in the magnetar wind 164. The detection of neutrinos would be crucial to breaking the degeneracies plaguing the electromagnetic data (which currently do not allow for a robust assessment of the source properties), allowing to pinpoint the mechanism powering the observed electromagnetic emission 162.
@@ 3.1.2 Neutron-star mergers
Following up the merger of two neutron stars, if the remnant compact object is a magnetar, its wind can interact with the ejecta, with protons being accelerated at the magnetosphere or the termination shock of the wind, see the right panel of Fig. 3. While particle acceleration is not efficient at the termination shock, neutrino emission from the magnetar could be detected in connection with gravitational waves 175; however, recent work suggests that the detection of such neutrinos may be more challenging than initially expected 176.
The non-detection of neutrinos from GRB 170817A 8, observed in connection with gravitational waves, is consistent with the fact that the relativistic jet was off-axis with respect to the observer 165; 64 (cf. Fig. 4 for an estimation of the muon neutrino fluence expected from the prompt emission). As the jet propagates within the ejecta, particle acceleration can take place at internal shocks inside the ejecta, leading to a neutrino signal within the detection reach of the IceCube Neutrino Telescope 54. Table 1 provides a summary of the energetics and the rate of the transients linked to the coalescence of neutron stars, with the related acceleration sites being summarized in the right panel of Fig. 3.
@@ 3.2 Best detection strategies for upcoming high-energy multi-messenger searches
The detection (or lack thereof) of high-energy neutrinos from transients stemming from collapsing massive stars or neutron-star mergers allows to gain crucial insight about the source properties, the behavior of plasma in highly magnetized media, as well as the efficiency of particle acceleration. Of course, the information provided by high-energy neutrinos should be complemented by the one carried by photons across wavebands and gravitational waves. Notably, detecting non-thermal neutrinos is especially crucial for electromagnetically dark sources, such as choked jets. In this case, we may expect a correlation of high-energy neutrinos with gravitational waves and $\mathcal{O}(10)$ MeV neutrinos (or interactions among such neutrinos 177) produced in the source core.
@@
Within the next decade, the number of electromagnetic detections of transients is expected to grow swiftly thanks to wide-field and high-cadence surveys, such as the soon-to-be operative Vera C. Rubin Observatory 178 and ULTRASAT 179, the Nancy Grace Roman Space Telescope 180 as well as the Cherenkov Telescope Array (CTA) 181. These surveys, among others, will complement the pioneering findings of the Zwicky Transient Facility (ZTF) 182, the Panoramic Survey Telescope and Rapid Response System 1 (Pan-STARRS1) 183, the Swift-Ultra-Violet/Optical Telescope 184, and the James Webb Space Telescope 185. At the same time, the next-generation gravitational-wave detectors, e.g. the Cosmic Explorer 186 and Einstein Telescope 187, are expected to largely expand the observational horizon explored by LIGO, Virgo, and KAGRA 188. For what concerns the detection of high-energy neutrino counterparts, the current observational landscape is dominated by the IceCube Neutrino Telescope 189 and the under-construction Cubic Kilometre Neutrino Telescope (KM3NeT) 190. The planned IceCube-Gen2 191, the Giant Radio Array for Neutrino Detection (GRAND200k) 192, and the Pacific Ocean Neutrino Experiment (P-ONE) 193, among several other neutrino detectors, are expected to strengthen and expand our detection opportunities (cf. Table 1 of 194 for an overview of existing and planned high-energy neutrino telescopes).
@@
To investigate the source properties and test our understanding of particle acceleration physics, it is however necessary to facilitate coordination as well as data exchange through multi-messenger programs. Several joint programs are already operative, for example ZTF, Pan-STARRS1, the Dark Energy Survey (DES), and the All-Sky Automated Survey for SuperNovae (ASAS-SN) carry out target-of-opportunity searches for optical counterparts of neutrino events and vice versa 195; 196; 197; 198; 199; similar collaborations exist for neutrino and gamma-ray joint searches, involving Fermi-LAT and the Imaging Atmospheric Cherenkov Telescopes (IACTs) 200; 201. To take full advantage of upcoming joint high-energy neutrino and electromagnetic detections, it is crucial to strengthen correlations between neutrinos, the radio waveband (likely carrying imprints of interactions with the circumstellar medium), and the X-ray one (sensitive to the activity of the central engine) to break the degeneracies intrinsic to the optical-ultraviolet-infrared band 51; 154; 163. The time window adopted to look for neutrinos should not be chosen arbitrarily; in fact, we expect that the neutrino signal can only be tightly linked to the waveband whose photons are produced in the same source region.
@@ 4 Neutrinos and explosive transients: future challenges and opportunities
The growing number of likely multi-messenger detections of transient astrophysical sources, as well as the major developments taking place in the field of neutrino telescopes and gravitational wave detectors, jointly with the upcoming wide-field and high-cadence surveys revolutionizing electromagnetic observations of transients, call for an improved understanding of the source physics and related particle production. Only by advancing the theoretical modeling, we would be able to take advantage of the upcoming plethora of data. On the one hand, it is essential to make progress in the modeling of neutrino flavor conversion in dense media. One challenge is that neutrino flavor conversion is not taken into account in (magneto)hydrodynamic simulations of the source. Hence, the impact of the feedback of flavor conversion physics on the source and the nucleosynthesis is yet to be understood. On the other hand, it is crucial to swiftly gain ground in our understanding of particle acceleration. Recent work highlights the importance of coupling particle transport consistently with general relativistic magnetohydrodynamics and particle-in-cell simulations to reliably assess the regions (and the efficiency) of particle acceleration and forecast the multi-messenger signals. For emerging source classes, neutrinos will be important to breaking the degeneracies affecting the electromagnetic signal, pinpointing the physics mechanisms powering such sources.
@@
The unprecedented number of real-time alerts expected to become available to telescopes involved in multi-messenger searches calls for a rapid development of the cyberinfrastructure needed to coordinate, cross-correlate, and analyze the large amount of data provided by such telescopes. In this respect, optimization of the communication infrastructure adopted to broadcast alerts is urgently needed. The Scalable Cyberinfrastructure to support Multi-Messenger Astrophysics (SCiMMA) 202 and the Time-domain Astronomy Coordination Hub (TACH) 203 are examples of projects in this direction.
Expanding the detection horizon beyond a single source, the diffuse emission of neutrinos from specific source classes can inform us on the source population properties. Besides complementing electromagnetic measurements, diffuse neutrino data could provide input on the sources that are dim or dark electromagnetically. However, to be able to extract details from the neutrino signal and predict the diffuse signal reliably, we should proceed on the modeling of stellar evolution, as well as better understand the role of a binary companion in the stellar evolution, and the cosmic rate.